Grade 11Algebra

Solve, 4x-5(2x)+4=0? plzzz assist me as quick as possible??

Surya
8 Years agoGrade 11

1 Answer

ayush yadav
ApprovedApproved Tutor Answer8 Years ago
22x-5(2x)+4=0
Now let 2x=a
Then
a2-5a+4=0
a2-4a-a+4=0
a(a-4)-1(a-4)=0
a=4
a=1
We had let 2x as a
2x=4 so x=2 
2x=1 so x=0
Its easy i hope it helps u